RC Update and Run plan for Aug 28,2020

From HallCWiki
Jump to: navigation, search

Back to the precious page

RC Meeting Information

RC Daily Meetings are at 4:15 PM, remote

General updates

  • 8/27 - Thursday: After the access ~ 1 hr data with reference cell (He3) (HMS is at 20 degrees -4 GeV and SHMS is at 18 degrees -5.6 GeV (Kin C-Z)
  • 8/27 - Thursday: 3:30 pm - 9:50pm data taking with PolHe3 transverse (HMS is at 20 degrees -4 GeV and SHMS is at 18 degrees -5.6 GeV (Kin C-Z)
  • Accelerator was down between 9:50 pm ( 8/27 - Thursday) - 1:50 am ( 8/28 - Friday):
    • RF trips (stormy weather)
  • 8/28 - Friday: 1:50am - 5:00 am - data taking with PolHe3 transverse (HMS is at 20 degrees -4 GeV and SHMS is at 18 degrees -5.6 GeV (Kin C-Z)
  • Accelerator has been down since 5:00 am:
    • SL RF went down because of iocch2a being unresolved this was followed by SL 2K cold box trip as it's unable to handle heat load changes with only half the SL up.
    • @11 am: reboot and restore iocch2a was successful and regain control over the SW linac signals. LHe overflow in SL return.
    • @1:49 pm: SL pump down is complete, RF recovery is ongoing: https://logbooks.jlab.org/entry/3836357
    • @2:49 pm: starting to recover RF https://logbooks.jlab.org/entry/3836382

Recent Hall Access (today 10:50 am - ?)

Recent NMR Measurements

Pending problem(s) / Things to track

  • Diagnose/resolve the transport ion chamber issue. (Investigation is ongoing by Jerry Kowal)
    • HV supply for transport ion chamber and Compton Ion chamber is split. Transport ion chamber is power cycled. We shall see the impact.
  • Spare reference cell options/status [Todd, JP]
  • Target mounting for Tommy (next cell) [Arun, et al]
  • Target mounting for Butterball [Arun, et al]
  • 1-pass changeover planning

Current Run Plan

  • Keep taking production data with PolHe3 transverse (HMS is at 20 degrees -4 GeV and SHMS is at 18 degrees -5.6 GeV (Kin C-Z))

List of Tasks for Down Time during Pass Change

  • Collimator installation (Confirmed with David H. (RadCon) today)
  • Purge water from compressor
  • Simona's SoLID Cherenkov work

General Reminders for Running

  • Make sure the raster size is correct (4--4.5mm)
  • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
  • Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
  • Always take 120Hz of EDTM
  • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
  • Change the IHWP at least every 8 hours when running on glass cells. Call Hall A in advance to request. If there are any issues, call the RC

Notices in Effect for Shift Crew

  • Follow all COVID rules and regulations as well as remind each other. Additional food guild lines:
    • All food are removed from the round table. Fruit are now store in the fridge, and individually packaged snacks are store on the shelf below the coffee machine.
    • No food near the SL and TO control station. If you would like to eat something, please use the empty round table (used to store food in the past), that is now empty and sanitized.
    • New coffee guide line: it is much preferred if you can bring you mug for the coffee. During the shift if one need to drink coffee near the control station, then the coffee mug must have a lid or have some kind of cover.
  • Beam positions: [1]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[2]
  • An new reference Cell refill procedure: New procedure
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • Slow ramping procedure of the SHMS, go small steps to the MOL, an great example (1 hour 30 min process)
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • New updates to the magnet momentum ramping procedure: https://hallcweb.jlab.org/wiki/index.php/How_to_operate_and_monitor_the_Spectrometer_Magnets
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instruction can be found: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT
  • New SHMS matrix element


Planned Accelerator Configuration Changes

  • Hall C to 1st pass on 8/31 and back to 5th pass on 9/03


  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920